[Brahms-l] Shift report 20040226 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Feb 26 2004 - 16:04:52 EST
------------------------------------------------------------
Date: Thu Feb 26 16:04:52 2004, Shift: 08:00-16:00
Supervisor: BB, with: YKL
Log book#15 pages 67 - 67

Shift Summary:
------------------------------------------------------------
1. Seven useful hours of beam time.

2. FS @8deg A1/2, MRS @4 deg A2255

   10605 begun by Eric Trig2 40K  Trig3 94K
   10606               Trig2 19K  Tirg3 72K
   10613                     59K        137K
   10614                     53K        199K
   10615                     19K         76K

3. Dana locked the key to FS so pectrometer cannot be moved
in. Flemming worked on general triggering, and removed
TRFS slats from trigger.


  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 26 16:05:16 2004

This archive was generated by hypermail 2.1.8 : Thu Feb 26 2004 - 16:05:42 EST